# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v2

@@ Abstract
In this work, we provide a self-contained derivation of the spin-operator matrix elements in the fermionic basis, for the critical periodic Ising chain at a generic system length $N\in 2Z_{\geq 2}$. The approach relies on the near-Cauchy property of certain matrices formed by the Toeplitz symbol in the critical model, and leads to simpler product formulas for the dressing functions in terms of square root functions. These square root products allow fully dis-singularized integral representations. In the finite volume scaling limit, they further reduce to the Binet’s second integral for the gamma function logarithm and its Hermite’s generalization. As such, all the matrix elements in the scaling limit allow simple product formulas in terms of the gamma function at integer and half-integer arguments, and are rational numbers up to $\sqrt{2}$. They are exactly the spin-operator form factors of the Ising CFT in the fermionic basis, whose explicit forms are much less well known in comparison to the finite-volume form factors in the massive theory. We also fully determine the normalization factor of the spin-operator and show explicitly how the coefficient $G(\frac{1}{2})G(\frac{3}{2})$ appear through a ground state overlap. As a by-product, we obtain a Mellin-Barnes representation of the Glaisher constant with two $\zeta$ functions
@@ 1 The critical Ising chain, square-root products and main results
The finite volume spin-operator form factors of the massive fermion Ising-QFT Fonseca:2001dc were well known, but what about their massless limits? Will the results become meaningless, or will they reduce to the spin-operator matrix elements of the Ising CFT in the fermionic basis, that were computed only iteratively in Yurov:1991my? To the author’s knowledge, it seems that many references who used the massive form factors in Fonseca:2001dc, including this paper itself, forgot to mention their convergence to a nontrivial finite volume massless scaling limit. Instead, the integrals in the leg functions were usually written in certain forms, creating impressions of their ill-posedness or triviality at the zero fermion mass [^1] For example, in Ref. Gabai:2019ryw, the leg function was written with an overall $m^{2}$. Still, it is easier to spot a non-trivial massless limit from the representation in Gabai:2019ryw than the rapidity space integral in Fonseca:2001dc..
@@
Similarly, in the literature, there are plenty discussions and derivations Lisovy:2001gc; Bugrij:2003xhk; Bugrij:2004sac; Gehlen_2008; Iorgov:2010yv; Iorgov:2010ie for the finite volume form factors in the Ising model and the Ising chain. These works are very non-trivial, but most of them were assuming $T>T_{c}$ or $T<T_{c}$, and forgot to mention what happened at the critical point. For example, in the published version of Gehlen_2008, the authors said explicitly after the Eq. (78) that the results are valid both when $g>1$ and $0\leq g<1$, but just omitted the critical point $g=1$. Very often, the formulas were written with a refactor such as $\xi=(g^{2}-1)^{\frac{1}{4}}$ in front, a factor that vanishes at the critical point $g=1$, like the expressions given to the leg functions in the massive theory, despite that in the proper way of taking the $g\rightarrow 1^{+}$ limit, the results are actually finite and as we will show, lead exactly to the form factors at the critical point.
@@
Meanwhile, in the literature, there were also not too many discussions on the spin-operator form factors in the Ising-CFT, despite that they are clearly universal, beautiful and unique quantities: the massless fermion basis of the Ising CFT is perhaps among the few integrable massless scattering basis Bazhanov:1994ft; Bazhanov:1996aq of 2D CFTs, that allow explicit expressions for local operator’s matrix elements between them. On the other hand, in comparison to the Virasoro basis, it could be the massless scattering basis, that respect better the scale-full deformations of CFTs either to the integrable lattice models in the UV Boos:2010qii, or to the massive QFTs in the IR. It is possible, that for certain non-inclusive high-energy limits that are more sensitive to the structures of the Hilbert space, one needs not just the standard CFT perturbation theory for point-like objects, but also the correct massless basis, to write out the correct high-energy asymptotics. These questions deserve more investigations, and were among the motivations of this work.
@@
Given all above, the author believes that it is reasonable and necessary to provide a self-contained derivation for the spin-operator form factors in the 2D periodic Ising chain, directly at the critical point. Below we will provide such a derivation, and show how in the finite volume scaling limit, the results converge to the Ising CFT versions. We believe that the products, integral representations and the large $N$ asymptotics in this work are simple yet beautiful. We also believe that the explicit formulas provided by this work could be of practical usages (such as the TFCSA applications Yurov:1991my).
@@ 1.2 The critical Ising chain
Here, $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{y}$ are the standard $2\times 2$ Pauli matrices at the site $i$. These spin operators commute at different sites. It is possible to choose the $\sigma^{x}$ and $\sigma^{z}$ to be purely real, as such $H$ can also be purely real. We define the standard Dirac matrices as Jordan:1928wi; Kaufman:1949ks
@@
To further set up our conventions for the eigensystem, here we review how the diagonalization process works in the $Z_{2}=-1$ (R) sector. The $Z_{2}=1$ (NS) sector can be handled similarly and is simpler. We will mainly follow the spirit of Kaufman:1949ks, but the computations are simpler in the Ising chain than the Ising model. For $Z_{2}=-1$, one has the following rotation matrix
@@
$c_{\beta}^{\dagger}=\frac{1}{2}\left(\gamma_{\beta}+i\gamma_{\beta+\frac{1}{2}}\right)\ ,\ c_{\beta}=\frac{1}{2}\left(\gamma_{\beta}-i\gamma_{\beta+\frac{1}{2}}\right)\ ,$ (30)
@@
Clearly, $H$ is invariant under $T_{+}$, while $H^{\prime}$ is invariant under $T_{-}$. Using the fermionization relations Eq. (9), one can show that
@@
Using the mode decompositions in Eq. (34) and Eq. (36), and the fact that $z^{N}=-1$, $\beta^{N}=1$, one has
@@
As such, when $\langle\Omega_{R}|\Omega_{NS}\rangle\neq 0$, due to the reality property, the phase difference $e^{-i(\theta_{+}-\theta_{-})}$ must be $\pm 1$ and is an absolute quantity independent of the state normalization. Latter we will show using the overlapping matrix $v_{z\beta}$ define in Eq. (128) that $e^{i(\theta_{+}-\theta_{-})}=1$.
@@ 1.3 Square-root products and form-factors at finite NN
$\langle\Omega_{R}|\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}|\Omega_{NS}\rangle\equiv(-1)^{l}{\cal M}_{2k+1,2l}(\beta_{i};z_{j})\ .$ (54)
@@
where $G({\cal Z})$ is the Barnes-$G$ function. The constant $(2\pi)^{\frac{1}{4}}G(\frac{1}{2})G(\frac{3}{2})$ will ensure that the rescaled corrector can be matched back to the infinite-volume version, at distances that are much larger than $1$, but much smaller than $N$. This is similar to the connecting property of the infinite-volume massive scaling functions Wu:1975mw. One major difference is: instead of appearing in the short distance expansion of the scaling function, the constant $G(\frac{1}{2})G(\frac{3}{2})$ moves to the ground state overlap, a quantity that resembles more the spontaneous magnetization in the infinite volume version. In fact, we will show that
@@
${\cal M}_{2k+1,2l}(\beta_{i};z_{j})=i^{k+l+1}\prod_{i<i^{\prime}}\frac{\sqrt{\beta_{i}}-\sqrt{\beta_{j}}}{\sqrt{\beta_{i}}\sqrt{\beta_{j}}-1}\prod_{i,j}\frac{\sqrt{\beta_{i}}\sqrt{z_{j}}-1}{\sqrt{\beta_{i}}-\sqrt{z_{j}}}\prod_{j<j^{\prime}}\frac{\sqrt{z_{j}}-\sqrt{z_{j^{\prime}}}}{\sqrt{z_{j}}\sqrt{z_{j^{\prime}}}-1}$
@@ 1.4 Scaling limits of the form factors
Clearly, this is because that when ${\cal Z}\rightarrow 1$, the integral in Eq. (76) receives the largest contribution only in the scaling region when $t-1={\cal O}(N^{-1})$. Here the crucial integral is
@@
This is a combination of the Binet’s second integral and its generalization by Hermite Hermite+1895+201+208; Boyack+2023. To convince oneself this non-trivial formula, one can use the following Barnes representation
@@
which allows to compute the small $\alpha$ expansion of $\mathfrak{b}(\alpha)$ and compare with that of the $\Gamma$ function ratios.
@@
and further label the non-zero quantization levels in the scaling limit to the left (L) and right (R) sector
$L:\beta_{n}^{+}=e^{i\frac{2\pi n}{N}},\ \beta_{n}^{-}=e^{-\frac{2\pi in}{N}}\ ,\ n\in Z_{\geq 1}\ ,$ (99)
$R:z_{m}^{+}=e^{i\frac{2\pi}{N}m},z_{m}^{-}=e^{-i\frac{2\pi}{N}m}\ ,\ m\in Z_{\geq 0}+\frac{1}{2}\ .$ (100)
@@
To present the results, we first clarify the conventions we used in the scaling limit. In all the matrix elements, we label the creation and annihilation operators from the left to the right. On each sides of the spin operator, the plus momenta (L) are always placed left to all the minus momenta (R). Without zero mode, we use the integer sequence $(n_{L},n_{R};m_{L},m_{R})$ in the subscript of the matrix element to denote that there are $n_{L}$ $\beta^{+}$s, $n_{R}$ $\beta^{-}$s, and $m_{L}$ $z^{+}$s, $m_{R}$ $z^{-}$s. And $(0,n_{L},n_{R};n_{L},n_{R})$ denotes that in addition there is a zero mode $c_{1}$, which is always placed next to the $\langle\Omega_{R}|$. The quantization levels for these modes are collected by four ordered sets $N_{L},N_{R},M_{L},M_{R}$, where $N_{L},N_{R}$ consist of positive integers, and $M_{L}$, $M_{R}$ consist of positive half integers. The ordering within each sets are the same as appearing in the matrix elements. We further introduce the following products
@@
where $N$ is a set positive integers, and $M$ is a set of positive half integers. A product equals to $1$ when one or both of the sets appearing in the product are empty. In terms of the products, the matrix elements without zero modes are given by
${\cal M}_{n_{L},n_{R};\ m_{L},m_{R}}\left(N_{L},N_{R};M_{L},M_{R}\right)=Z_{n_{L},n_{R};m_{L},m_{R}}D(N_{L})D(N_{R})D(M_{L})D(M_{R})$
@@
For the two particle functions involving $\beta=1$, simply notice that in our ordering they all become 1-1. Collecting all the 1-1 and the $e^{\pm\frac{i\pi}{4}}$ factors leads to the final expressions Eq. (106) for the phase factors. The explicit expressions Eq. (104), Eq. (105) and Eq. (106) for the form factors in the scaling limit are the second major result of the paper. They provide explicit formulas for the CFT matrix elements computed recursively in Yurov:1991my. Notice that all the form factors in the scaling limit are rational numbers in our convention, due to the overall $e^{\pm\frac{i\pi}{4}}$ in the phase factors.
@@ 2 Derivation of the main results
After introducing all the main results, in this section we derive the results presented in the previous section. The overall strategy is similar to cmp/1103901557; cmp/1103904079, but the way of solving the matrix equation is different. Rather than using the Wiener-Hopf trick suitable for the infinite-volume limit, our approach is based on the near-Cauchy property of the $W$, $W^{\prime}$ matrices given below in Eq. (125), which is a special property of the critical model and allows explicit expressions at finite $N$.
@@
Since both of the $|\Omega_{R}\rangle$ and $|\Omega_{NS}\rangle$ are gaussian-type ground states of fermionic quadratic forms defined in terms of the same set of fermionic operator, multi-point correlators between the two states are all generated by the overlapping matrices through Wick-contractions Perk:1983hg. In particular, since the $\sigma^{x}_{0}=\Gamma_{0}$ is also a fermionic operator, correlators with $\sigma$ can also be obtained from Wick contractions. It turns out that it is more convenient to introduce the following matrices
@@
Knowing the conversion rules, we now consider the following identity cmp/1103901557
@@
At this stage, it is time to introduce the following matrices
@@
and move the $c_{\beta}^{\dagger}$s in $c_{\bar{z}}$ to the left, one can show that all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}c_{\beta_{4}}|\Omega_{NS}\rangle$ are also zero. Recursively [^4] This method can also be used to prove the Wick theorem in the current context, as it leads to a recursive relation for Pfaffians cmp/1103901557., this implies that all the matrix elements $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ are zero. But this is impossible, because $c_{\beta_{1}}^{\dagger}c^{\dagger}_{\beta_{2}}...c^{\dagger}_{\beta_{n}}|\Omega_{R}\rangle$ is a complete set of orthogonal basis of the full $2^{N}$ dimensional Hilbert space, and vanishing of all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ (odd numbers naturally vanish due to the charge condition) would imply that $|\Omega_{NS}\rangle\equiv 0$. We will show below that $W$ is indeed invertible. As such, by dividing the $\langle\Omega_{R}|\Omega_{NS}\rangle$ on the two sides of Eq. (124) one obtains the matrix equation

### v3

@@ Abstract
In this work, we provide a self-contained derivation of the spin-operator matrix elements in the fermionic basis, for the critical periodic Ising chain at a generic system length $N\in 2Z_{\geq 2}$. The approach relies on the near-Cauchy property of certain matrices formed by the Toeplitz symbols in the critical model, and leads to simpler product formulas for the dressing functions in terms of square root functions. These square root products allow fully dis-singularized integral representations. In the finite volume scaling limit, they further reduce to the Binet’s second integral for the gamma function logarithm and its Hermite’s generalization. As such, all the matrix elements in the scaling limit allow simple product formulas in terms of the gamma function at integer and half-integer arguments, and are rational numbers up to $\sqrt{2}$. They are exactly the spin-operator form factors of the Ising CFT in the fermionic basis, whose explicit forms are much less well known in comparison to the finite-volume form factors in the massive theory. We also fully determine the normalization factor of the spin-operator and show explicitly how the coefficient $G(\frac{1}{2})G(\frac{3}{2})$ appear through a ground state overlap. As a by-product, we obtain a Mellin-Barnes representation of the Glaisher constant with two $\zeta$ functions
@@ 1 The critical Ising chain, square-root products and main results
The finite volume spin-operator form factors of the massive fermion Ising-QFT Fonseca and Zamolodchikov (2001) were well known, but what about their massless limits? Will the results become meaningless, or will they reduce to the spin-operator matrix elements of the Ising CFT in the fermionic basis, that were computed only iteratively in Yurov and Zamolodchikov (1991)? To the author’s knowledge, it seems that many references who used the massive form factors in Fonseca and Zamolodchikov (2001), including this paper itself, forgot to mention their convergence to a nontrivial finite volume massless scaling limit. Instead, the integrals in the leg functions were usually written in certain forms, creating impressions of their ill-posedness or triviality at the zero fermion mass [^1] For example, in Ref. Gabai and Yin (2022), the leg function was written with an overall $m^{2}$. Still, it is easier to spot a non-trivial massless limit from the representation in Gabai and Yin (2022) than the rapidity space integral in Fonseca and Zamolodchikov (2001)..
@@
Similarly, in the literature, there are plenty discussions and derivations Lisovy (2002); Bugrij and Lisovyy (2003, 2004); Gehlen et al. (2008); Iorgov et al. (2011); Iorgov and Lisovyy (2011) for the finite volume form factors in the Ising model and the Ising chain. These works are very non-trivial, but most of them were assuming $T>T_{c}$ or $T<T_{c}$, and forgot to mention what happened at the critical point. For example, in the published version of Gehlen et al. (2008), the authors said explicitly after the Eq. (78) that the results are valid both when $g>1$ and $0\leq g<1$, but just omitted the critical point $g=1$. Very often, the formulas were written with a refactor such as $\xi=(g^{2}-1)^{\frac{1}{4}}$ in front, a factor that vanishes at the critical point $g=1$, like the expressions given to the leg functions in the massive theory, despite that in the proper way of taking the $g\rightarrow 1^{+}$ limit, the results are actually finite and as we will show, lead exactly to the form factors at the critical point.
@@
Meanwhile, in the literature, there were also not too many discussions on the spin-operator form factors in the Ising-CFT, despite that they are clearly universal, beautiful and unique quantities: the massless fermion basis of the Ising CFT is perhaps among the few integrable massless scattering basis Bazhanov et al. (1996, 1997) of 2D CFTs, that allow explicit expressions for local operator’s matrix elements between them. On the other hand, in comparison to the Virasoro basis, it could be the massless scattering basis, that respect better the scale-full deformations of CFTs either to the integrable lattice models in the UV Boos et al. (2010), or to the massive QFTs in the IR. It is possible, that for certain non-inclusive high-energy limits that are more sensitive to the structures of the Hilbert space, one needs not just the standard CFT perturbation theory for point-like objects, but also the correct massless basis, to write out the correct high-energy asymptotics. These questions deserve more investigations, and were among the motivations of this work.
@@
Given all above, the author believes that it is reasonable and necessary to provide a self-contained derivation for the spin-operator form factors in the 2D periodic Ising chain, directly at the critical point. Below we will provide such a derivation, and show how in the finite volume scaling limit, the results converge to the Ising CFT versions. We believe that the products, integral representations and the large $N$ asymptotics in this work are simple yet beautiful. We also believe that the explicit formulas provided by this work could be of practical usages (such as the TFCSA applications Yurov and Zamolodchikov (1991)).
@@ 1.2 The critical Ising chain
Here, $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{y}$ are the standard $2\times 2$ Pauli matrices at the site $i$. These spin operators commute at different sites. It is possible to choose the $\sigma^{x}$ and $\sigma^{z}$ to be purely real, as such $H$ can also be purely real. We define the standard Dirac matrices as Jordan and Wigner (1928); Kaufman (1949)
@@
To further set up our conventions for the eigensystem, here we review how the diagonalization process works in the $Z_{2}=-1$ (R) sector. The $Z_{2}=1$ (NS) sector can be handled similarly and is simpler. We will mainly follow the spirit of Kaufman (1949), but the computations are simpler in the Ising chain than the Ising model. For $Z_{2}=-1$, one has the following rotation matrix
@@
$c_{\beta}^{\dagger}=\frac{1}{2}\left(\gamma_{\beta}-i\gamma_{\beta+\frac{1}{2}}\right)\ ,\ c_{\beta}=\frac{1}{2}\left(\gamma_{\beta}+i\gamma_{\beta+\frac{1}{2}}\right)\ ,$ (30)
@@
Clearly, $H$ is invariant under $T_{+}$, while $H^{\prime}$ is invariant under $T_{-}$. Using the fermionization relations Eq. (9) and Eq. (45), one can show that
@@
Using the mode decompositions in Eq. (34) and Eq. (36), and the fact that $z^{N}=-1$, $\beta^{N}=1$, one has Iorgov and Lisovyy (2011)
@@
As such, when $\langle\Omega_{R}|\Omega_{NS}\rangle\neq 0$, due to the reality property, the phase difference $e^{-i(\theta_{+}-\theta_{-})}$ must be $\pm 1$ and is an absolute quantity independent of the state normalization. Latter we will show explicitly using the overlapping matrix $v_{z\beta}$ define in Eq. (131) that $e^{i(\theta_{+}-\theta_{-})}=1$. Equivalently, the left continuity at $h=1$ and the absence of discontinuities when $h\in[0,1)$ allows $e^{i(\theta_{+}-\theta_{-})}$ to be computed by the $h=0$ version, which is clearly $1$. Indeed, when $h=0$, $H=H^{\prime}$, and the absolute ground state is a trivial “all-down state”.
@@ 1.3 Square-root products and form-factors at finite NN
$\langle\Omega_{R}|\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}|\Omega_{NS}\rangle\equiv(-1)^{l}{\cal M}_{2k+1,2l}(\{\beta_{i}\};\{z_{j}\})\ .$ (54)
@@
where $G({\cal Z})$ is the Barnes-$G$ function. The constant $(2\pi)^{\frac{1}{4}}G(\frac{1}{2})G(\frac{3}{2})$ will ensure that the rescaled corrector can be matched back to the infinite-volume version, at distances that are much larger than $1$, but much smaller than $N$. This is similar to the connecting property of the infinite-volume massive scaling functions Wu et al. (1976). One major difference is: instead of appearing in the short distance expansion of the scaling function, the constant $G(\frac{1}{2})G(\frac{3}{2})$ moves to the ground state overlap, a quantity that resembles more the spontaneous magnetization in the infinite volume version. In fact, we will show that
@@
${\cal M}_{2k+1,2l}(\{\beta_{i}\};\{z_{j}\})=i^{k+l+1}\prod_{i<i^{\prime}}\frac{\sqrt{\beta_{i}}-\sqrt{\beta_{j}}}{\sqrt{\beta_{i}}\sqrt{\beta_{j}}-1}\prod_{i,j}\frac{\sqrt{\beta_{i}}\sqrt{z_{j}}-1}{\sqrt{\beta_{i}}-\sqrt{z_{j}}}\prod_{j<j^{\prime}}\frac{\sqrt{z_{j}}-\sqrt{z_{j^{\prime}}}}{\sqrt{z_{j}}\sqrt{z_{j^{\prime}}}-1}$
@@ 1.4 Scaling limits of the form factors
Clearly, this is because that when ${\cal Z}-1={\cal O}(N^{-1})$, the integral in Eq. (76) receives the largest contribution only in the scaling region when $t-1={\cal O}(N^{-1})$. Here the crucial integral is
@@
This is a combination of the Binet’s second integral and its generalization by Hermite Hermite (1895); Boyack (2023). To convince oneself this non-trivial formula, one can use the following Barnes representation
@@
which allows to compute the small $\alpha$ expansion of $\mathfrak{b}(\alpha)$ and compare with that of the $\Gamma$ function ratios. Notice that throughout this work, the contours for such Barnes integrals are always along the imaginary direction from ${\rm Re}(u)-i\infty$ to ${\rm Re}(u)+i\infty$.
@@
and further seperate the non-zero quantization levels in the scaling limit to the plus ($L$) and minus ($R$) sectors
$L:\beta_{n}^{+}=e^{i\frac{2\pi n}{N}};\ R:\beta_{n}^{-}=e^{-i\frac{2\pi n}{N}}\ ,\ n\in Z_{\geq 1}\ ,$ (99)
$L:z_{m}^{+}=e^{i\frac{2\pi m}{N}};\ R:z_{m}^{-}=e^{-i\frac{2\pi m}{N}}\ ,\ m\in Z_{\geq 0}+\frac{1}{2}\ .$ (100)
@@
To present the results, we first clarify the conventions we used in the scaling limit. In all the matrix elements, we label the creation and annihilation operators from the left to the right. On each sides of the spin operator, the plus momenta (L) are always placed left to all the minus momenta (R). Without zero mode, we use the integer sequence $(n_{L},n_{R};m_{L},m_{R})$ in the subscript of the matrix element to denote that there are $n_{L}$ $\beta^{+}$s, $n_{R}$ $\beta^{-}$s, and $m_{L}$ $z^{+}$s, $m_{R}$ $z^{-}$s. And $(0,n_{L},n_{R};n_{L},n_{R})$ denotes that in addition there is a zero mode $c_{1}$, which is always placed next to the $\langle\Omega_{R}|$. The quantization levels for these modes are collected by four ordered sets $N_{L},N_{R},M_{L},M_{R}$, where $N_{L},N_{R}$ consist of positive integers, and $M_{L}$, $M_{R}$ consist of positive half integers. To summarize, in the scaling limit, the momenta in the matrix elements are collected by the following ordered sets
$\{\beta_{i}\}_{(N_{L},N_{R})}=\{\beta_{n}^{+},n\in N_{L}\}\bigcup\{\beta_{n}^{-},n\in N_{R}\}\ ;\ |N_{L}|=n_{L},\ |N_{R}|=n_{R}\ ,$ (103)
$\{z_{j}\}_{(M_{L},M_{R})}=\{z_{m}^{+},m\in M_{L}\}\bigcup\{z_{m}^{-},m\in M_{R}\}\ ;\ |M_{L}|=m_{L},\ |M_{R}|=m_{R}\ ,$ (104)
$\{\beta_{i}\}_{(0,N_{L},N_{R})}=\{1\}\bigcup\{\beta_{i}\}_{(N_{L},N_{R})}\ ,$ (105)
where the orderings are always from the left to the right, and $|A|$ denotes the size of the set $A$. To present the scaling limit, we further introduce the following products
@@
where $N$ is a set positive integers, and $M$ is a set of positive half integers. A product equals to $1$ when one or both of the sets appearing in the product are empty. In terms of the products, scaling limits of the matrix elements without zero mode are given by
${\cal M}_{n_{L}+n_{R};\ m_{L}+m_{R}}\left(N_{L},N_{R};M_{L},M_{R}\right)=Z_{n_{L},n_{R};m_{L},m_{R}}D(N_{L})D(N_{R})D(M_{L})D(M_{R})$
@@
For the two particle functions involving $\beta=1$, simply notice that in our ordering they all become 1-1. Collecting all the 1-1 and the $e^{\pm\frac{i\pi}{4}}$ factors leads to the final expressions Eq. (109) for the phase factors. The explicit expressions Eq. (107), Eq. (108) and Eq. (109) for the form factors in the scaling limit are the second major result of the paper. They provide explicit formulas for the CFT matrix elements computed recursively in Yurov and Zamolodchikov (1991). Notice that all the form factors in the scaling limit are rational numbers in our convention, due to the overall $e^{\pm\frac{i\pi}{4}}$ in the phase factors.
@@ 2 Derivation of the main results
After introducing all the main results, in this section we derive the results presented in the previous section. The overall strategy is similar to Abraham (1978a, b) [^4] This strategy is essentially equivalent to that of Iorgov and Lisovyy (2011), but allows to avoid the explicit Gaussian wave functions of the ground states. , but the way of solving the matrix equation is different. Rather than using the Wiener-Hopf trick suitable for the infinite-volume limit, our approach is based on the near-Cauchy property of the $W$, $W^{\prime}$ matrices given below in Eq. (128), which is a special property of the critical model and allows explicit expressions at finite $N$.
@@
Since both of the $|\Omega_{R}\rangle$ and $|\Omega_{NS}\rangle$ are gaussian-type ground states of fermionic quadratic forms defined in terms of the same set of fermionic operator, multi-point correlators between the two states are all generated by the overlapping matrices through Wick-contractions Perk et al. (1984). In particular, since the $\sigma^{x}_{0}=\Gamma_{0}$ is also a fermionic operator, correlators with $\sigma$ can also be obtained from Wick contractions. It turns out that it is more convenient to introduce the following matrices
@@
Knowing the conversion rules, we now consider the following identity Abraham (1978a)
@@
At this stage, it is time to introduce the following $N\times N$ matrices
@@
and move the $c_{\beta}^{\dagger}$s in $c_{\bar{z}}$ to the left, one can show that all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}c_{\beta_{4}}|\Omega_{NS}\rangle$ are also zero. Recursively [^5] This method can also be used to prove the Wick theorem in the current context, as it leads to a recursive relation for Pfaffians Abraham (1978a)., this implies that all the matrix elements $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ are zero. But this is impossible, because $c_{\beta_{1}}^{\dagger}c^{\dagger}_{\beta_{2}}...c^{\dagger}_{\beta_{n}}|\Omega_{R}\rangle$ is a complete set of orthogonal basis of the full $2^{N}$ dimensional Hilbert space, and vanishing of all the $\langle\Omega_{R}|c_{\beta_{1}}c_{\beta_{2}}c_{\beta_{3}}...c_{\beta_{2n}}|\Omega_{NS}\rangle$ (odd numbers naturally vanish due to the charge condition) would imply that $|\Omega_{NS}\rangle\equiv 0$. We will show below that $W$ is indeed invertible. As such, by dividing the $\langle\Omega_{R}|\Omega_{NS}\rangle$ on the two sides of Eq. (127) one obtains the matrix equation